in need of help, solve the system:
{x^2-100y^2=100
{x^2+ y^2=100.
thanks for your help!!!

Answer by stanbon(75887)

{x^2-100y^2=100
{x^2+ y^2=100.
-------------
Subtract the 1st equation from the 2nd to get:
101y^2 = 0
y^2 = 0
y = 0
----------
Substitute into x^2 + y^2 = 100 to solve for "x":
x^2 + 02 = 100
x = +10 or x= -10
=====================
Solution: (10,0) and (-10,0)
===============
Cheers,
Stan H.